Hello, i want to ask you Inventory cost =3.6(B1+B2)/2+3.75(B2+B3)/2+3.98(B3+B4)/2+4.28(B4+B5)/2+4.2(B5+B6)/2+3.9(B6+B7)/2, So, how to compute 3.6, 3.75, 3.98, 4.28, 4.2? And why 3.6X(B1+B2) and divide by 2? Also, i want to ask the objective function for this problem why is minimize, not maximize? Thank you for your help!The Upton Corporation manufactures heavy-duty air compressors for the home and light industrial markets. Upton is presently tThe basic decision Uptons management team faces is how many units to manufacture in each of the next six months. We will repConstraints 2,000 <P < 4,000 1,750 <P2 < 3,500 2,000 <P3 < 4,000 2,250 <P4 < 4,500 2,000 <P5 < 4,000 1,750 <P < 3,500 product
Answer #1

1. How to calculate inventory carrying cost of 3.6, 3.75, 3.98, 4.28, 4.20, 3.90..

Solution : These are inventory carrying cost that is equal to 1.5% of production cost of unit in same month ( refer paragh above the table in question)

Month 1 : 240 *1.5%= 3.6

Month 2 : 250*1.5% = 3.75

Month 3 : 265 *1.5% =3.98

Month 4: 285* 1.5% = 4.28

Month 5 = 280*1.5% = 4.20

Month 6 =260*1.5% = 3.90

Question 2. Why to divide by 2?

Solution : Since the inventory that needs to be considered for calculating the value of inventory carrying cost is to be average of inventory at the beginning and end of month.. the opening and closing inventory has been added and divided by 2 to reach at the average inventory..

Question 3. Why the objective is to minimise?

Solution : Objective is to minimise the Total inventory cost, both manufacturing and carrying cost. This is done to reduce expense and have a higher profit margins.